A project with a life of 5 has an initial fixed asset investment of $24,360, an initial NWC investment of $2,320, and an annual OCF of –$37,120. The fixed asset is fully depreciated over the life of the project and has no salvage value.

  

If the required return is 8 percent, what is the project's equivalent annual cost, or EAC?

A project with a life of 5 has an initial fixed asset investment of $24,360, an initial NWC investment of $2,320, and an annual OCF of –$37,120. The fixed asset is fully depreciated over the life of the project and has no salvage value.
  
If the required return is 8 percent, what is the project's equivalent annual cost, or EAC?
  • $-36,895.71

  • $-34,662.09

  • $-43,406.72

  • $-45,577.06

  • $-41,236.38
Answer #1
CF0 -26680
CF1 -37120
CF2 -37120
CF3 -37120
CF4 -37120
CF5 -34800
NPV -$173,310.44
EAC -$43,406.72

CF0 = -24,360 - 2,320 = -26,680 and CF5 = -37,120 + 2,320 = -34,800

Using the NPV function in excel or calculator, we can calculate the present value of these cash flows with 8% discount rate.

NPV = -173,310,44

Now, EAC can be calculated using PMT function

N = 5, PV = -173,310.44, FV = 0, I/Y = 8% => Compute PMT = $43,406.72